### v3

@@ 7 Robustness and sensitivity checks
In this section, we conduct a number of robustness checks to support our main conclusions. First, we examine whether tax optimization, rather than organizational costs, drives the location choice of intermediate jurisdictions. In the first column of Table 7, we introduce controls for corporate tax differentials, calculated as the ratio of the average profit tax in the destination to the origin, using data from the World Bank’s Doing Business (2019). The profit tax measures the total burden on a typical firm as a percentage of commercial profits. These additional results confirm that investment shares in the $ikj$ triplet increase when country $k$ offers a relatively more favorable tax environment. However, while corporate tax differentials play a role, they do not diminish the importance of organizational costs in shaping ownership chains.
A second concern is whether labor costs, which are particularly important for multinational firms that coordinate global supply chains, are the most important determinant of the choice of the intermediate jurisdiction. In column 2, we extend our baseline model by incorporating labor cost differentials, measured as the ratio of employee compensation in the destination to the origin, using data from the World Bank’s Doing Business (2019). Not surprisingly, we find a negative and significant relationship between labor costs and investment decisions, but only in the bilateral relationship between countries $k$ and $i$. Higher labor costs in country $k$ discourage investment through that country, but the main findings on overlapping working hours remain significant, with estimated coefficients consistent in magnitude with the baseline results.
@@
$\sqrt{\frac{\hat{\mathcal{C}}_{ij}}{K}}$135.109*** -135.218*** -141.157***
@@
  • Note: Ratios $CT_{k}/CT_{i}$ and $CT_{j}/CT_{k}$ control for relative advantage in corporate taxation. Ratios $LC_{k}/LC_{i}$ and $LC_{j}/LC_{k}$ control for the relative advantage in labor costs. Given the presence of zeros in the case of direct control, we scale the distances in km by using the hyperbolic sine transformation ($ln(x+\sqrt{x^{2}+1})$), which allows approximating to the natural logarithm while retaining the zeros (Bellemare and Wichman, 2020). Standard errors are double-clustered by origin and destination in parentheses (*** p$<$0.01, ** p$<$0.05, * p$<$0.1).
@@
$\sqrt{\frac{\hat{\mathcal{C}}_{ij}}{N_{K}}}$89.573*** -158.517*** -118.742***
@@
  • Note: Sectors are defined according to the NAICS 2017 classification at the 2-digit level. Given the presence of zeros in the case of direct control, we scale the distances in km by using the hyperbolic sine transformation ($ln(x+\sqrt{x^{2}+1})$), which allows approximating to the natural logarithm while retaining the zeros (Bellemare and Wichman, 2020). Standard errors are double-clustered by origin and destination in parentheses (*** p$<$0.01, ** p$<$0.05, * p$<$0.1).
@@ 8 Conclusions
This study examines how multinational enterprises (MNEs) structure complex hierarchical ownership networks, where control is often exercised through multi-tiered chains of subsidiaries spanning multiple national borders. Although these structures play a crucial role in the global organization of MNEs, they have been largely overlooked in the literature, likely due to data limitations. Using a unique global database, we first document stylized facts on ownership chains and then develop a theoretical framework to explain their geography and structure.
Our findings reveal that most affiliates are controlled through ownership chains and that complex multinational enterprises with extensive corporate structures play a dominant economic role. In the upper tail of the distribution, large corporate networks control more than 100 subsidiaries, operate in more than 19 countries, and span more than 15 industries.
Beyond their economic significance, we examine the geographic organization of complex ownership chains, finding that ownership chains can span up to seven national borders and that their complexity increases with firm size. In particular, we find a correlation between the hierarchical position of subsidiaries and their geographic distance from headquarters. Subsidiaries positioned further down the hierarchy tend to be located farther from the parent company. Furthermore, using overlapping working hours across time zones as a proxy for the ease of managerial communication, we find that subsidiaries at the bottom of the hierarchy share fewer working hours with managers at the top.
@@
Our results highlight how little is still known about the ownership structures of multinational firms. While official statistics often fail to fully capture global ownership networks, the increasing standardization of national business registries is improving data availability. Looking ahead, we see promising opportunities for future research on the determinants and consequences of multinational ownership chains, a field that remains ripe for deeper exploration.
@@
Appendix B: Naïve empirical models of parent-subsidiary relationships
In the following paragraphs, we introduce three naïve unstructured applications of gravity models to observe correlations, which are informative for the setup of our theoretical model, and to discuss what the analyst misses when not considering information on ownership chains. We argue that this section is important to identify empirical problems from previous literature. In each case, we test the following general Pseudo-Poisson specification at the country level:
${N_{ij}=exp(\bm{\beta}\mathbf{X}_{ij}+\gamma_{i}+\gamma_{j})\epsilon_{ij}}$ (10)
where $N_{ij}$ counts the number of companies in country $j$ controlled by companies in country $i$. We include bilateral frictions in matrix $\mathbf{X}_{ij}$. Notably, besides classical gravity controls sourced from the CEPII Gravity dataset (Conte et al., 2022) $wh_{ij}$, we include a measure of overlapping working hours between countries $i$ and $j$ as a proxy of the common working schedule across time zones. See further details in Section 3. Finally, $\gamma_{i}$ and $\gamma_{j}$ capture origin and destination country fixed effects. [^22] Please note that our information on MNEs’ perimeters is cross-sectional. Therefore, we can focus on the between-variation and cannot observe the within-variation.
At this point, we are ready with the empirical model, but how do we treat ownership chains? We will show that each of the three following cases is wrong. Or, we could say that they are reduced forms of the real model. Let us introduce the issue starting from a fictional sample, including two kinds of MNEs most common in our data set, whose fictional structures are represented in Figure B1.
Figure B1: A fictional dataset
  • Note: This fictional dataset includes a simple and a complex MNE. Each node represents a company, and each link represents a company’s controlling voting rights in another company.
Figure B2: Implementations of naïve bilateral gravity
  • Notes: The figure shows three fictional cases of ownership chains included in bilateral gravity models. The legends on the right report the (directed) links one would include in each application.
More specifically, in Figure B1, we draw one simple MNE controlling only one subsidiary, which we know is the most common case, although we also know that, on aggregate, these MNES have a small economic weight (see Figures 3 and 4 ). A second complex MNE on the right controls seven subsidiaries organized along four ownership chains: i) A $\rightarrow$ B; ii) A $\rightarrow$ C $\rightarrow$ E $\rightarrow$ H; iii) A $\rightarrow$ D $\rightarrow$ F; iv) A $\rightarrow$ D $\rightarrow$ G.
Now, how would an analyst include this sample in a bilateral gravity model?
Let us first assume that linkages are not disclosed, but the set of subsidiaries is known. In this case, the analyst will count the parents in country $i$ and the subsidiaries in country $j$. This is the case (a) in the fictional sample redrawn in Figure B2. We guess this is the case that most scholars have in mind when testing gravity models with firm-level data on MNEs. See, for example, Mutti and Grubert (2004), which reports an empirical analysis using data on the location choices of US parent companies.
A second alternative is that ownership chains are only partially disclosed because the statistician has problems collecting them. S/he falsely attributes the ultimate control of a subsidiary to its immediately controlling corporate shareholder. Hence, our fictional complex MNE will appear fragmented into three smaller pieces. The information included in Eq. 10 is false. This is the second case in Figure B2, where the list of links on the right does not include $C\rightarrow E$. Indeed, full disclosure of the corporate structure of MNEs is a long-standing problem for national and international statistics offices. The necessity to identify ultimate controlling entities is required by several guidelines (OECD, 2005; UNCTAD, 2009; IFRS, 2011), seldom complied with.
A third alternative is that the analyst knows the entire ownership chain but does not know how to fit in a bilateral gravity. S/he might consider every pairwise link connecting two firms. As from Figure B2, it would imply an inflation of linkages. The analyst breaks the unity of ownership linkages, while every middlemen subsidiary is technically considered as a parent company originating its own investment choice. Please note that this setup reflects the common approach of using FDI data to proxy MNE relationships. See Bénassy-Quéré et al. (2007); Stein and Daude (2007); Egger and Pfaffermayr (2004); Blonigen et al. (2003). FDI data provided by the OECD is collected following the 4th edition of the Benchmark Definition of FDI (BMD4), which primarily attributes inward FDI stocks to the immediate partner country in the transaction.
In Table B1, we report the results of our three naïve applications. The main coefficients are consistent with the magnitude and sign we would expect from these frictions, except in case a), where the coefficient on the N. of overlapping working hours is not statistically significant. In either case, findings might be irrelevant because we know that each application misses a point, and they are wrong overall because they neglect the existence of ownership chains, which are an essential ingredient in the organization of MNEs.
Table B1: Naïve bilateral gravities
Dep. var. $N_{ij}$= # of companies country i controls in country j
Sample: Case (a) Case (b) Case (c)
N. of overlapping 0.015 0.056** 0.058**
working hours (0.023) (0.024) (0.023)
(0.186) (0.150) (0.154)
(log of) distance -0.340*** -0.264*** -0.252***
(0.077) (0.090) (0.088)
Home 2.377*** 3.324*** 3.355***
(0.295) (0.245) (0.251)
Language 0.741*** 0.772*** 0.780***
(0.090) (0.104) (0.102)
Colony dependence 0.312** 0.326** 0.304**
(0.124) (0.132) (0.130)
Legal origins 0.289*** 0.373*** 0.377***
(0.089) (0.082) (0.084)
RTA 0.188 0.319** 0.292*
Observations 39,585 39,786 39,786
Fixed effects i,j i,j i,j
  • Note: The figure shows results from three applications of naïve gravity models, where the information on ownership chains is absent in Case (a), it is only partially reconstructed in Case (b), or it is considered but breaking the ownership chains in bilateral segments in Case (c). Standard errors clustered by origin country in parentheses (*** p$<$0.01, ** p$<$0.05, * p$<$0.1).
Appendix C: Tables and graphs
Table C1: Country-level covariates and sources
@@
Table C2: Sample coverage by industry
@@
Table C3: Sample coverage by size
@@
Figure C1: Financial accounts - geographic coverage in 2019
@@
Figure C2: Size distribution - (log of) number of controlled subsidiaries
@@
Table C4: Lengths of ownership chains

### v4

@@ 7 Robustness and sensitivity checks
In this section, we conduct several robustness checks to support our main conclusions. First, we examine whether tax optimization, rather than organizational costs, drives the location choice of intermediate jurisdictions. In the first column of Table 7, we introduce controls for corporate tax differentials, calculated as the ratio of the average profit tax in the destination to the origin, using data from WorldBank (2019). The profit tax measures the total burden on a typical firm as a percentage of commercial profits, and this is the same variable that has been used for preliminary evidence in Section 3. These additional results confirm that investment shares in the ikjikj triplet increase when country $k$ offers a relatively more favorable tax environment. However, while corporate tax differentials play a role, they do not diminish the importance of organizational costs in shaping ownership chains.
A second concern is whether labor costs, which are particularly important for multinational firms that coordinate global supply chains, are the most important determinant of the choice of the intermediate jurisdiction. In column 2, we extend our baseline model by incorporating labor cost differentials, measured as the ratio of employee compensation in the destination to that in the origin, using data from WorldBank (2019). Not surprisingly, we find a negative and significant relationship between labor costs and investment decisions, but only in the bilateral relationship between countries $k$ and $i$. Higher labor costs in country $k$ discourage investment through that country, but the main findings on overlapping working hours remain significant, with estimated coefficients consistent in magnitude with the baseline results.
@@
$\sqrt{\frac{\hat{\mathcal{C}}_{ij}}{K}}$ -135.109*** -135.218*** -141.157***
@@
  • Note: Ratios $CT_{k}/CT_{i}$ and $CT_{j}/CT_{k}$ control for relative advantage in corporate taxation. Ratios $LC_{k}/LC_{i}$ and $LC_{j}/LC_{k}$ control for the relative advantage in labor costs. Given the presence of zeros in the case of direct control, we scale the distances in km by using the hyperbolic sine transformation ($ln(x+\sqrt{x^{2}+1})$), which allows approximating to the natural logarithm while retaining the zeros (Bellemare and Wichman, 2020). Standard errors are double-clustered by origin and destination in parentheses (*** p<<0.01, ** p<<0.05, * p<<0.1).
@@
$\sqrt{\frac{\hat{\mathcal{C}}_{ij}}{N_{K}}}$ -89.573*** -158.517*** -118.742***
@@
  • Note: Sectors are defined according to the NAICS 2017 classification at the 2-digit level. Given the presence of zeros in the case of direct control, we scale the distances in km by using the hyperbolic sine transformation ($ln(x+\sqrt{x^{2}+1})$), which allows approximating to the natural logarithm while retaining the zeros (Bellemare and Wichman, 2020). Standard errors are double-clustered by origin and destination in parentheses (*** p<<0.01, ** p<<0.05, * p<<0.1).
@@ 8 Conclusions
This study examines how multinational enterprises (MNEs) structure complex hierarchical ownership networks, where control is often exercised through multi-tiered chains of subsidiaries spanning multiple national borders. Although these structures play a crucial role in the global organization of MNEs, they have been largely overlooked in the literature, likely due to data limitations. Using a unique global database, we first document stylized facts on the geography of ownership chains and then develop a theoretical framework based on the idea that communication frictions among managers play a crucial role.
When we examine the geographic organization of complex ownership chains, we find they can be long and strech accross multiple country borders. Approximately 61% of subsidiaries are positioned at least two ownership links’ distance from the parent company, 47% cross at least two country borders, and 53% extend approximately more than 10,000 kilometers.
In particular, we find that there are two relevant correlations: i) between the hierarchical position of subsidiaries and their geographic distance from headquarters; ii) between the hierarchical position of subsidiaries and the difference in time zones, i.e, the overlapping in working hours. The latter is for us a measure of the ease of communication among managers across countries. Briefly, subsidiaries at the bottom of the hierarchy share fewer working hours with managers at the top.
@@
Our results highlight the limited understanding of the ownership structures of multinational firms, despite their significant weight in the global economy. While official statistics often fail to fully capture global ownership networks, the increasing standardization of national business registries is improving data availability and accuracy. Looking ahead, we see promising opportunities for future research on the emergence and dynamics of multinational ownership chains.
@@
Appendix B: Tables and graphs
Table B1: Country-level covariates and sources
@@
Total Tax Contribution Rate (TTCR)
Total amount of taxes and mandatory contributions payable by businesses, as a percentage
of commercial profit.
WorldBank (2019)
@@
Table B2: Sample coverage by industry
@@
Table B3: Sample coverage by size
@@
Figure B1: Financial accounts - geographic coverage in 2019
@@
Figure B2: Size distribution - (log of) number of controlled subsidiaries
@@
Table B4: Lengths of ownership chains
@@
Appendix C: Additional stylized facts on multinational enterprises and their corporate structures
A key observation is that most subsidiaries of MNEs are controlled indirectly through ownership chains. Table C1 classifies all subsidiaries in our dataset based on the type of control exercised by the parent companies, regardless of geographic location or firm size. We find that 54% of subsidiaries are indirectly controlled, underscoring the prevalence of multi-tiered corporate structures. Referring to the hypothetical ownership chain in Figure 2, directly controlled subsidiaries maintain a direct link to the parent company, while indirectly controlled subsidiaries are linked through other subsidiaries, forming multilevel ownership chains.
Table C1: Direct and indirect control of subsidiaries
Type of control N. subsidiaries %
Direct control 827,516 45.81
Indirect control 978,952 54.19
Total 1,806,468 100.00
Complex multinational enterprises that operate through ownership chains account for a disproportionate share of global economic activity, accounting for 95% of total MNE revenues, despite representing only 27% of the firms in our dataset. In Figure C1, we distinguish between MNEs that exclusively use direct control over subsidiaries (left panel) and MNEs that include at least one ownership chain (right panel). While the latter group is numerically smaller, it dominates global sales, reinforcing the idea that larger corporate networks tend to be structured through multi-tiered ownership chains. [^21] Our findings are consistent with previous studies by Rungi et al. (2017); Altomonte and Rungi (2013), which show that the largest percentile of MNEs by corporate perimeter accounts for approximately 75% of global sales. This pattern emerges because MNEs without ownership chains tend to be smaller in size. On average, they generate 0.107 billion USD in revenue, compared with 1.782 billion USD for complex MNEs with chains of ownership. [^22] Ownership data tend to be more comprehensive than financial accounts because tax authorities require full disclosure of shareholding structures, while financial reporting requirements vary across jurisdictions. As a result, financial data may be incomplete due to exemptions and reporting discrepancies. In our analysis, we assume that missing sales data are not systematically correlated with firm size. Alternative imputation methods do not significantly alter our results, as confirmed in Table C1. Appendix Figure B1 illustrates that sample selection bias is often driven by geographic coverage rather than firm size.
Figure C1: MNEs, ownership chains, and their economic weight
MNEs exhibit considerable heterogeneity in their organizational structure, varying in the number of subsidiaries, ownership chains, host countries, and industries in which they operate. In Figure C2, we classify MNEs according to the number of subsidiaries they control (x-axis) and report their frequency distribution (y-axis). The first bar represents the simplest MNEs, which control only one subsidiary. More complex MNEs, owning at least two subsidiaries, can choose to structure them along an ownership chain. The distribution is highly skewed, with a long right tail, indicating that only 1.2% of MNEs control more than 100 subsidiaries. This heterogeneity goes beyond the number of subsidiaries controlled. MNEs also vary significantly in the number of countries in which they operate and the diversity of industries in which their subsidiaries operate, further highlighting the complexity of global corporate structures.
Figure C2: MNEs’ corporate perimeters
Table C2 provides further evidence of the heterogeneity of MNEs’ corporate structures. Looking first at the number of subsidiaries controlled, we find an average of 7.93 subsidiaries per parent company. But this masks a highly skewed distribution. The median MNE controls only one subsidiary, while the third quartile of MNEs reaches three subsidiaries per parent company. Moving to the right tail, we find that MNEs at the 90th percentile have more than 10 subsidiaries, and only at the 99th percentile do we reach the most complex MNEs, which control more than 120 subsidiaries, with a maximum of 5,075 subsidiaries worldwide. [^23] This is the case of Berkshire Hathaway, the largest multinational conglomerate holding company in the United States, ultimately owned by Warren Buffett. [^24] In Appendix Figure B2, we estimate the distribution of controlled subsidiaries using a negative binomial regression, which yields a mean of 1.882 and an overdispersion parameter of 0.581. The figure visually represents a log-log plot of the frequency distribution of subsidiaries per parent. A similarly skewed pattern is observed when focusing on intermediate affiliates. The median MNE has zero intermediate affiliate, but the right tail expands sharply, with a jump from 2 to 20 subsidiaries between the 90th and 99th percentiles.
Looking at geographic dispersion, Table C2 shows that the average MNE operates in 2.81 countries. Since our definition of an MNE requires at least one foreign subsidiary, the minimum number of countries is two. However, the 90th percentile includes firms operating in four countries, while at the 99th percentile the number jumps to 19 countries and reaches a maximum of 147.252525This extreme case corresponds to DHL, the globally recognized logistics and courier group. Finally, industry diversification follows a similar pattern. The most complex MNEs, those in the right tail of the distribution, operate in more than 15 industries at the 99th percentile, reaching a maximum of 64 industries. [^26] According to our data, this is the case of HSBC Holdings PLC, originally founded as the Hongkong and Shanghai Banking Corporation and now a British banking group headquartered in London.
Table C2: Heterogeneous distributions of MNEs
Variable Mean st. dev. Min p50 p75 p90 p99 Max
N. controlled subsidiaries 7.93 50.09 1 1 3 10 120 5,075
N. middlemen 1.18 7.72 0 0 1 2 20 688
N. of industries 2.19 3.01 1 1 2 4 15 64
N. of countries 2.81 3.92 2 2 2 4 19 147
The relationship between the number of subsidiaries controlled by an MNE and the maximum length of its ownership chains is illustrated in Figure C3. The plot represents the predicted ranges of a regression model in which the dependent variable is the logarithm of the firm’s scope and the maximum chain length (also in logs) is the key explanatory variable. The relationship appears to be non-linear, with an estimated elasticity of 2.01, suggesting that ownership chains tend to lengthen disproportionately as firms grow. However, large MNEs do not always adopt deep hierarchical structures. Some firms maintain flat ownership networks in which all subsidiaries remain directly controlled by the parent company. [^27] For example, Assyce Fotovoltaica SL, a Spanish provider of solar energy equipment, directly controls approximately 300 subsidiaries without intermediate affiliates.
Figure C3: Number of controlled subsidiaries and maximum ownership chain length